### v1

# Elements of World Knowledge (EWoK): A cognition-inspired framework for evaluating basic world knowledge in language models

Abstract
The ability to build and leverage world models is essential for a general-purpose AI agent. Testing such capabilities is hard, in part because the building blocks of world models are ill-defined. We present Elements of World Knowledge (EWoK), a framework for evaluating world modeling in language models by testing their ability to use knowledge of a concept to match a target text with a plausible/implausible context. EWoK targets specific concepts from multiple knowledge domains known to be vital for world modeling in humans. Domains range from social interactions (help/hinder) to spatial relations (left/right). Both, contexts and targets are minimal pairs. Objects, agents, and locations in the items can be flexibly filled in enabling easy generation of multiple controlled datasets. We then introduce EWoK-core-1.0, a dataset of 4,374 items covering world knowledge domains. We evaluate open-weights large language models (B–B parameters) across a battery of evaluation paradigms along with a human norming study comprising 12,480 measurements. The overall performance of all tested models is worse than human performance, with results varying drastically across domains. These data highlight simple cases where even large models fail and present rich avenues for targeted research on LLM world modeling capabilities.
1 Introduction
Large language models (LLMs) acquire a substantial amount of knowledge from their training data, both through direct memorization and through learning co-occurrence-based text patterns (Bender and Koller, 2020; Grand et al., 2022; Pavlick, 2022). This knowledge comprises both knowledge about language (e.g. word meanings and rules of syntax) and knowledge about the world (e.g. social conventions and physical properties of objects). Contemporary LLMs’ knowledge of language is very robust (Mahowald et al., 2024), as evidenced by the fluency of the text that they generate. But how robust is their understanding of the basic social, physical, and relational concepts that are foundational to our everyday experience?
In this paper, we present Elements of World Knowledge (EWoK)111Data and associated code are available at: http://ewok-core.github.io, a flexible framework for evaluating world modeling in LLMs (see Figure 1 for an overview). The EWoK framework consists of: (a) several domains that constitute the foundation for basic human world knowledge and are processed by dedicated cognitive and neural systems; (b) specific concepts within each domain; (c) a set of item templates (and modular components to procedurally generate these templates) that test knowledge of these concepts by contrasting plausible and implausible context–target combinations; (d) a set of fillers to populate the templates, such that each template can be used multiple times; (e) a pipeline to generate a specific set of items (a dataset) based on these source materials.
Why elements?
Our framework targets specific cognitive concepts (or concept pairs, such as friend/enemy). Concept knowledge is not limited to definitions, but can be used productively across a wide range of scenarios. Thus concepts leveraged in context are the first-class object of the EWoK framework, as opposed to individual sentences or facts. This approach stands in contrast to many NLP benchmarks, which often aim to evaluate knowledge based on individual items, or to develop sophisticated tasks that require drawing on many concepts simultaneously. Although real-world tasks can be more useful for comparing model performance, individual item complexity makes it hard to assess why a model fails. For instance, LLMs exhibit mixed performance on theory of mind tasks, which has led to disagreements about whether LLMs fail at inferring mental states or auxiliary cognitive abilities, such as knowing what it means for a box to be “transparent” (Kosinski, 2023; Ullman, 2023). Our framework mitigates this problem by explicitly linking the items with the concepts that they test.
Why cognition-inspired?
World knowledge is a notoriously fuzzy capability; which knowledge domains should one focus on? Here, we select a range of domains that have been shown to recruit dedicated cognitive and/or neural machinery in humans, such as knowledge of intuitive physics (McCloskey, 1983; Battaglia et al., 2013), knowledge of physical and spatial relations (Hafri and Firestone, 2021), intuitive number sense (Dehaene, 2011), social reasoning (Carlson et al., 2013), and reasoning about agents that involves both physical and social knowledge (Liu et al., 2024). These knowledge domains are not specific to language (Jackendoff, 2002); in fact, many are present in preverbal infants (Spelke and Kinzler, 2007). However, language contains a rich amount of information that reflects grounded world knowledge (Roads and Love, 2020; Abdou et al., 2021; Patel and Pavlick, 2021), such that LLMs might in principle acquire such domain-specific knowledge from text alone.
Why plausibility?
To evaluate basic world modeling, we use combinations of plausible vs. implausible context–target pairs. Plausibility here serves as a proxy for factual accuracy: instead of deciding whether a factual statement is true or false, a model needs to determine whether a given scenario makes sense (is plausible). Having an accurate world model is necessary for consistently distinguishing plausible and implausible scenarios no matter how they are worded.
Why minimal pairs (of pairs)?
Both contexts and targets in EWoK have a minimal-pairs design, such that a specific targeted change to a sentence (e.g., left right) results in an opposite result (plausible/implausible combination). This approach can help identify specific manipulations that LLMs are and are not sensitive to, with the goal of targeted diagnostics. It can also be leveraged in the future for mechanistic interpretability research.
Why context–target combinations?
LLMs have a remarkable capacity for memorization, such that many plausible and implausible sentences can be distinguished solely based on their presence in the training data (e.g., The fox chased the rabbit is more common than The rabbit chased the fox). In contrast, our framework tests LLMs’ ability to evaluate contextual plausibility, such that the same exact target (The piano is left of Ali.) is either plausible or implausible depending on the context (see Figure 1 right).
Our paper is structured as follows. In Section 2, we review prior research related to this topic. In Section 3, we describe the core components of the EWoK framework. In Sections 4 and 5, we describe our evaluation strategy and the practical steps that we take to reduce the risk of dataset contamination. In Section 6, we show LLM performance on EWoK-core-1.0, a dataset generated via the EWoK framework. In Section 7, we discuss our results and future prospects for this line of work.
2 Related Work
The capabilities we evaluate are closely related to the notion of commonsense knowledge, an area that has numerous text-based benchmarks (e.g., Levesque et al., 2012; Sakaguchi et al., 2021; Zellers et al., 2019), including targeted evaluations of physical (Bisk et al., 2020) and social commonsense (Sap et al., 2019). LLMs often struggle on such commonsense benchmarks, likely due to the reporting bias in their training data (Shwartz and Choi, 2020): conversations and texts typically do not include commonly observed or obvious information (Gordon and Van Durme, 2013). Thus, although our items target basic world knowledge and are not designed to be challenging, LLMs might still struggle with them due to the reporting bias.
One specific version of reporting bias affects perceptually grounded knowledge. Co-occurrence information that is easily available through perception (e.g., the fact that bananas are typically yellow or wheels are typically round) is often underrepresented in language corpora. This bias has led an earlier generation of language models to underperform on physically and perceptually grounded world knowledge tasks (Lucy and Gauthier, 2017; Utsumi, 2020) and exhibit representational differences in physical features that are more/less talked about (Abdou et al., 2021; Lewis et al., 2019). That said, in spite of such biases, models trained on text do learn a substantial amount of distributional information from perceptual domains (Roads and Love, 2020; Abdou et al., 2021; Sorscher et al., 2022), meaning that much of the perceptual world knowledge can be acquired without grounding. Overall, we expect LLMs to perform above-chance on world knowledge domains that are perceptually grounded (such as physical relations or material properties), although they might be more challenging than, e.g., social domains.
Our evaluation is also related to works on natural language inference and entailment. The recognizing textual entailment (RTE) task Dagan et al. (2010) poses two sentences (a text expression and and hypothesis ) to a system and asks it to determine whether follows from . The natural language inference (NLI) task follows a similar challenge (Bowman et al., 2015; Williams et al., 2018; Conneau et al., 2018) and involves making a 3-way judgment about whether a premise entails, contradicts, or is neutral relative to a hypothesis. EWoK asks whether a target sentence is plausible given a context , which might—but does not have to—indicate an entailment relationship between the two. Though widely successful as a challenge, a large body of subsequent work has highlighted issues with RTE- and NLI-style evaluation: language models can often use heuristics (such as artifacts left behind by human annotators and lexical statistics) to “solve” the task without meaningful semantic understanding or reasoning (Poliak et al., 2018; Liu et al., 2020; McCoy et al., 2020; Gururangan et al., 2018). We address this limitation by (1) posing the task as a minimal pair where each of two targets must be preferred over its alternative given the right context, making reliance on target plausibility alone impossible; (2) annotating minimal pair contrast type to test whether which item design features drive model performance; and (3) testing the relationship between LLM performance and surface-level item properties, such as item length, average word frequency, and performance of a baseline bag-of-words embedding model.
Our approach to dataset design is similar in spirit to the bAbi framework (Weston et al., 2016), which used simple synthetic tasks probing world knowledge and reasoning; however, our items are both simpler in design (they target individual concepts and do not, as of yet, require multi-chain reasoning or transitive inference) and harder in practice (a minimal pair, context-dependent design greatly reduces the availability of response heuristics, a serious problem in bAbi; Kaushik and Lipton, 2018).
The minimal pair design is common in datasets inspired by psycholinguistics and cognitive science, such as SyntaxGym (Gauthier et al., 2020), BLiMP (Warstadt et al., 2020), and COMPS (Misra et al., 2023). In particular, it has previously been used to evaluate the models’ ability to distinguish plausible and implausible events (Pedinotti et al., 2021; Kauf et al., 2023), a task that draws heavily on commonsense knowledge. The popular Winograd Schema Challenge also had a minimal pairs setup (Levesque et al., 2012), even though that requirement was relaxed in later versions. We here extend this approach by employing a minimal pairs-of-pairs design, where both context and target sentences have a minimal pair counterpart.
Until 2023, the dominant approach of assessing language model performance on a minimal pair has been to calculate each item’s (pseudo) log probability under the model. This method is effective at distinguishing grammatical and ungrammatical sentences (e.g., Warstadt et al., 2020), plausible and implausible events (Kauf et al., 2023), and relevant vs. irrelevant object properties (Misra et al., 2023), while often being calibrated to human sentence ratings and processing costs (Lipkin et al., 2023; Shain et al., 2024). Yet raw log probabilities reflect a number of surface-level properties of the input, such as word frequency (Kauf et al., 2023) and the number of possible paraphrases (Holtzman et al., 2021). An alternative approach, recently made possible with more powerful LLMs, is to prompt an LLM to to rate item plausibility, either absolute (on a Likert scale) or relative to the other item in the minimal pair. This approach can theoretically result in more task-specific estimates. However, for a range of linguistic and word prediction tasks, LLMs actually perform worse with direct prompting than via implicit log probability assessment (Hu and Levy, 2023), likely because of additional task demands imposed by the need to decipher instructions in the prompt (Hu and Frank, 2024). Thus, we report both log probability comparisons and explicit prompting results.
3 The Framework
We provide a flexible generative synthetic data pipeline (Figure 1), capable of producing many diverse datasets, each with unique specifications and statistics, while preserving metadata and decision traces. In Section 6, we use this framework to generate EWoK-core-1.0, a systematic, broad-coverage, context-sensitive world knowledge dataset containing 4,400 items.
Item format
Effective use of world knowledge incorporates both access to a rich set of priors about the world’s structure and the ability to integrate this information on-the-fly with surrounding context. Thus, our items challenge models to leverage concepts in context. Each item consists of two minimal pair contexts (e.g., : The piano is in front of Ali. Ali turns left., : The piano is in front of Ali. Ali turns right) and two minimal pair target sentences (e.g., : The piano is right of Ali., : The piano is left of Ali.). The two target concepts are juxtaposed such that in any item, and . Thus, base target probabilities and cannot serve as plausibility cues: a model has to rely on context to establish plausibility.
Domains & concepts
EWoK is designed around domains of general world knowledge, with the current set shown in Figure 3. Each domain includes a set of concepts. For example, the domain social relations includes friend, enemy, teacher, student, boss and subordinate, and so on.
Dataset generation procedure
Each concept is associated with several items that test knowledge of the concept (often, but not always, by contrasting it with another concept). Items are created in a flexible yet controlled manner using the EWoK dataset generation procedure. At the core of the framework, we have atomic units and combination rules that, subject to constraints, lead to the generation of templates which are then populated with fillers—this leads to generating many more carefully-controlled items than a single-step template-filling approach could generate.
Contexts & targets
A target is a simple sentence that incorporates a concept; each domain has one or more associated targets depending on the lexical properties of the concepts in that domain. A contrasting target pair is generated using one of two mechanisms: concept swap, which contrasts the same target with different concepts filled in, and variable swap, which swaps two objects or agents mentioned in the target (only possible for certain concepts). For instance, {agent1} is to the left of {agent2} can be contrasted with {agent1} is to the right of {agent2} (concept swap) or with {agent2} is to the left of {agent1} (variable swap).
A context pair is one or more minimal pair(s) of sentences that can each be paired with a target pair such that but not matches and but not matches . They are typically specific to an opposing concept pair (left/right) or to a single concept (left; only possible when variable swap is allowed). Similarly, a contrasting context pair is generated using one of two mechanisms, namely, a filler swap and a variable swap. A filler swap uses contrasting fillers to generate two versions of the item. A variable swap changes positions of two entities of the same kind to change the nature of the relationship or dynamic between them as long as the concept supports it.
Templates & fillers
Each collection of concepts, contexts, and targets can then be compiled to a set of templates, partial items with typed variables describing the range of fillers for which the intended meaning holds. For example, in the template {object2:can_bounce=True} bounced off {object1} from below, any object may fill {object1}, e.g., the desk or the crate, whereas {object2:can_bounce=True} must be filled by an object marked with a flag for can_bounce=True, e.g., the ball or the tire. Across domains we list over filler items across classes with type restrictions.
When populating the templates with fillers, users can specify various custom parameters, e.g.: (1) specify the number of items to generate from each template; each full set of items is referred to as a version; (2) select whether fillers should be held constant across all items in a version or allowed to vary; and (3) apply transformations to filler restrictions at compile-time, for instance restricting all agents to take non-Western names via agent->agent:western=False, or swapping all objects with nonce words via object->nonword. Such flexibility allows for controlled experimentation of the features modulating model performance. With these arguments and others easily accessible from the command line interface, users are supported to generate an enormous space of EWoK variations, and may easily add new domains, concepts, fillers, and more.
4 Evaluation
Using the EWoK framework, we compiled the EWoK-core-1.0 dataset by generating unique fixed substitutions of filler items across templates from domains. In this section, we describe the evaluation of EWoK-core-1.0 via three different evaluation paradigms: traditional plausibility estimates via querying the probability of sentences under the model, LogProbs, as well as two prompt-based strategies, Likert and Choice. The majority of the results reported use the LogProbs evaluation method, which allows comparing a wide range of base and finetuned models. As we show, LogProbs outperforms direct prompting even for large and/or instruction-tuned models.
For the prompt-based evaluations, we collected data from both LLMs as well as human participants using paired identical prompts. Drawing inspiration from comparative psychology, such an approach of matched evaluation has been proposed as a way to support increasingly fair evaluations of LLMs, allowing for more direct comparison of performance, consistency, and context-sensitivity with human participants (Lampinen, 2022).
4.1 Scoring metrics
For LogProbs evaluation, we use token-level LLM probabilities to calculate as a sum of conditional log probabilities of each token: , where is the vector of tokens composing the target . For Likert, participants (humans and models) are prompted to rate the plausibility of the concatenation of each and pair on a 1–5 scale. For Choice, participants (humans and models) are presented with and , followed by a single target ( or ), and then prompted to select the context ( or ) that better matches the target. For Likert and Choice, details about text generation hyperparameters can be found in A.3 and exact prompt templates can be found in Appendix A.4. The metric for correctness of a given item is the recovery of the designed item structure that and , where score reflects for LogProbs, the integer rating for Likert, and correct context index selection for Choice. In all cases, participants must correctly identify both matches to get the full score ( point). Identification of only one match receives points. In the case of the Likert task, if a model returns the same rating for both pairs, the model receives points. Such a paradigm supports a trivial baseline for all scenarios. Either a random coin flip or a deterministic model generating the same response for each query independent of context will trivially achieve this baseline.
4.2 Models
We evaluated =20 transformer language models, selected to span a few points in the model design space. Models primarily vary in size (# of parameters; ranging from B–B) and pre-training diet (both # of tokens and source of training corpora). While most evaluated LLMs are dense pre-trained transformers (=13), there are a few one-off comparisons supported including the presence of supervised fine-tuning for instructions (=4) or chat (=2), and mixture-of-experts (MoE) ensembling (=1). We do not intend to draw conclusions about any of these design decisions, but rather to expose variation. Aside from these considerations in exploring variation, the selection of fine-tuned models was filtered to those that did not require specific formatting via the use of a prompt template. Since we evaluate LLMs and humans on identical prompts, it was critical to have complete flexibility in formatting. The full set of evaluated models are listed in Fig. 4 as well as in Appendix A.2.
4.3 Surface-level item properties
We additionally tested a baseline bag-of-words model based on word2vec embeddings (Mikolov et al., 2013). Embeddings for each word in a context or a target were summed together to derive one vector per context/target. The context/target match was determined by a cosine similarity metric, such that an item is scored correctly if > (and vice versa for ). We then examined the extent to which LLM and human performance correlates with the word2vec baseline. We also tested whether LLM performance correlates with the number of words in an item, as well as with average word frequency in an item using unigram counts from the Google Ngrams (Michel et al., 2011) 2012 American English corpus.
4.4 Human data
For independent norming of the items in EWoK-core-1.0, we collected data from human participants. We recruited a total of =1,262 participants (591 female, 579 male, 27 other; median age 36; all US-residents with English as their first language) via Prolific, an online study platform. Raters with poor agreement with others () were excluded. The task participants performed was nearly identical to the Likert version of the task for LLMs (in a pilot study, we determined that human results for Likert and Choice show correlation; see Appendix A.1 & B.1 for more details).
5 Release Considerations
Our release-related goals are to (a) reduce the chances of accidental incorporation of EWoK items into LLMs’ training data and (b) promote accountability and reporting when such incorporation is done intentionally. Thus:
EWoK-core-1.0 is released on HuggingFace Datasets (Lhoest et al., 2021) with gated user access to prevent scrapers from accessing it automatically. Users will simply accept a CC-BY license and accompanying Terms of Use (TOU) wherein they will agree to explicitly report any instances when a language model was trained on the EWoK-core-1.0 items, and will be granted access automatically. Link: https://huggingface.co/datasets/ewok-core/EWoK-core-1.0
The code for the EWoK item generation framework is shared in a separate repository on GitHub, with template files downloadable as password-protected archives to prevent automatic scraping. The repository is also protected with a TOU that requires anyone training or fine-tuning on any data generated using EWoK to report that fact. Link: https://github.com/ewok-core/ewok
The code required to replicate the results in this paper, along with human study and model performance data, is shared as a separate GitHub repository following the same protections. Link: https://github.com/ewok-core/ewok-paper.
Although not flawless, this strategy allows us to minimize the chances of accidental incorporation of our data into training sets, as well as to send a clear signal to model developers about the importance of accountable reporting of evaluation dataset use during model training.
6 Experiments: EWoK-core-1.0
EWoK-core-1.0 is challenging for LLMs
Despite the fact that EWoK-core-1.0 was not designed to be a hard dataset, we found that even the larger models generally perform much below humans: the best model tested, falcon-40b-instruct, has an accuracy of whereas humans perform at (Table AB.2). As expected, larger models tend to do better, although the performance gains are modest. Instruction tuning does not consistently increase or decrease LLM performance under the LogProbs metric.
Performance varies drastically by domain
Figure 4 shows model results by domain, from easiest to hardest (see also Table AB.2). Social interactions is the easiest for both LLMs (mean=, best=) and humans (); spatial interactions is both hardest for LLMs (mean=, best=) and shows the largest gap with human performance (who score highly at ). Overall, domain difficulty is consistent across LLMs, such that hard domains are hard for all and vice versa. One notable exception is the heterogeneous performance of the phi models, with phi-1 consistently among the worst models, phi-1.5 outperforming all models and even humans on physical dynamics, and phi-2 ranging from on par with the largest models on some domains to worse than gpt2-xl on spatial relations. These results can perhaps be attributed to their unique training procedure, which focuses prominently on LLM-generated synthetic data.
LLMs show heterogeneous performance across dataset versions
Our generation framework is designed to allow easy substitution of different values for names, objects, and locations, among other variables. In principle, these values should not affect the results ( still best matches and not ). However, models show somewhat different performance on the 5 different versions of the dataset, with phi-2 and phi-1.5 showing the largest performance range ( for both). This heterogeneity indicates that arbitrary item choices can substantially affect model performance, such that, in our dataset and in others, “1%” improvements might not be truly meaningful because they would not generalize. Humans show somewhat heterogeneous performance too (range ), although that heterogeneity is driven only by a subset of the domains (Figure 4).
Domain content, item design features, and surface-level item features all affect LLM performance
It is critically important to test whether cross-domain performance differences are indeed driven by domain content or are rather attributable to other factors. We consider item design features (direct/indirect context; contrast type for context sentences; and contrast type for target sentences) and surface-level features (average word frequency and sentence length). Figure 5 shows that all these factors affect model performance, often in different ways than they affect humans. Performance of a bag-of-words word2vec-based baseline on different dataset domains and versions is predictive of LLM but not human performance (although baseline performance overall is only ). The number of words in an item negatively affects LLM but not model performance. Unexpectedly, we found a negative relationship between word frequency and both LLM and human performance; follow-up examination of the data showed that this effect is driven by the fact that physical-relations and spatial-relations, the two hardest domains for LLMs, have the highest word frequency. To evaluate relative contributions of these factors to model performance, we jointly modeled all features using mixed effects regression. Design features and surface-level features all contributed to model performance (Table AB.2), with word frequency having a significant positive effect and the number of words having a significant negative effect, as expected. Importantly, domain remained a significant predictor of performance even when accounting for other factors.
LogProbs yield higher accuracy than prompting
Evaluating LLMs with LogProbs resulted in above-chance accuracy for almost all models (Figure 6). As expected, the gap between LogProbs and prompting was larger for smaller models (Hu and Frank, 2024). We used the same prompting setup across all models to assess them in the same way as humans (Lampinen, 2022) (although note that models are also aided via -shot examples). It is possible that targeted prompt-engineering or alternative evaluation strategies will result in above-LogProbs performance, a direction we leave to future work.
Human ratings are often, but not always accurate
Finally, we examined discrepancies between human ratings and experimental labels. Sometimes, the discrepancy resulted from specific fillers changing the plausibility of a pair; for instance The cooler is inside the car. Chao cannot see the cooler. is implausible because the cooler is large and the car has windows, although, for smaller objects and containers without windows, the scenario is plausible. Sometimes, humans made mistakes. One such example is cardinal directions from the spatial relations domain. The scenario The bakery is north of Chao. Chao turns around. The bakery is south of Chao. is implausible because cardinal directions do not depend on the agent’s orientation, and yet our participants often marked it as plausible. Overall, human data collection is a valuable source of information on our dataset but it does not replace ground truth labels.
7 Discussion
We present a systematic, flexible framework that can be used to test basic world knowledge in language models. Our goal was to develop a dataset that: (1) uses a uniform item format to probe diverse domains of physical and social knowledge, (2) contains items that probe specific concepts (“elements of world knowledge”) within these domains, (3) requires integrating information across sentences, such that the same target sentence is plausible given one context and implausible given another, and (4) consists of generic templates that can be used to generate a wide variety of items.
We then presented evaluation results for a set of openly available models on EWoK-core-1.0, a dataset generated using the EWoK framework. This dataset is moderately challenging for LLMs, with performance varying substantially across domains. We also show that LogProbs contain enough information for most LLMs to perform above chance and that evaluation via prompting underperforms relative to LogProbs even for better-performing models, in agreement with prior results (Hu and Frank, 2024; Hu and Levy, 2023; Hu et al., 2024; Kauf et al., 2024).
The EWoK framework opens up multiple avenues for future work:
Targeted experiments
The flexibility of our framework allows conducting specific experiments using customized sets of fillers. For instance, one might investigate whether LLMs perform differently on items that include western vs. nonwestern names, items that refer to people by names vs. longer descriptors (“the man in the black hat”), or even items featuring nonwords (like “florp”) instead of real object names.
Interpretability research
Knowledge editing research (e.g., Meng et al., 2022, 2023) has often focused on encyclopedic knowledge; but what about knowledge of basic physical and social concepts? Our controlled minimal pair stimuli can allow researchers to identify and manipulate model circuits that might be selectively responsible for knowledge of specific concepts from an array of domains.
From elements to world models
For a model to function as a flexible and robust general purpose AI system, it needs to be able to construct, maintain, and update internal world models (Ha and Schmidhuber, 2018; LeCun, 2022) (in cognitive science, variants of such world models are also known as mental models or situation models). The extent to which LLMs possess and use internal world models is subject to ongoing investigation (Hao et al., 2023; Yildirim and Paul, 2024; Wong et al., 2023). The EWoK framework offers an opportunity to combine individual elements of world knowledge to construct multi-step scenarios for evaluating world modeling capabilities in LLMs, within and across physical and social knowledge domains.
Limitations
Our dataset is written in English; LLM performance might be lower on other languages, especially under-resourced ones. Adapting the EWoK framework to other languages might require redesigning the set of concepts and materials we use, which are currently grounded in the English lexicon. Thus, a multilingual framework can help more cleanly dissociate linguistic and conceptual effects on model performance. Another limitation is that we use the same prompting setup for all models. With tailored prompt engineering, alternative generation methods, or chain-of-thought, LLM performance could improve. Regardless, we expect model generation to conform with general world knowledge under a wide range of evaluation scenarios—the fact that they do not is informative. Finally, due to the synthetic nature of our dataset, some items are semantically weird. Although not necessarily a problem, an alternative is to use LLMs to populate templates. We leave this approach to future work noting that it can introduce confounds that inflate model performance (e.g., Panickssery et al., 2024).
8 Conclusion
To evaluate the ability of LLMs to construct robust world models, we need to test their ability to reason about the fundamental elements of world knowledge. The EWoK framework provides a way to systematically evaluate such knowledge, highlights that LLMs continue to fall short on simple scenarios requiring physical, spatial, or social knowledge, and offers opportunities for further targeted evaluations of LLMs.
Acknowledgements
We thank everyone who contributed ideas on dataset design, especially Hayley Ross and Yuhan Zhang. We thank members of LINGO Lab at MIT for providing feedback on earlier versions of the work. This work was supported by the Language Mission of the MIT Quest for Intelligence.
Appendix A Methods
A.1 Human data collection
We collected human data in two phases: a pilot study used to validate our task on a single domain and determine the measurement technique to use for data collection (Choice or Likert), and a main study where we repeated data collection for the full set of EWoK-core-1.0 items. In the pilot study we determined that human judgments are measurement-technique-invariant (Ivanova et al., 2024), so we did not collect Choice judgments on the full set of materials, instead relying on Likert judgments to score items. Results from both studies are discussed in Appendix B.1.
Pilot study
The pilot study was done on materials from one of the EWoK-core-1.0 subdomains—social relations—using one set of variables to populate the fillers (not used in the main study). Participants from USA were recruited using Prolific, an online study platform, based on being self-reported native and fluent English speakers. We recruited a total of 30 participants across conditions. Of these, 18 reported identifying as ‘female’, 11 as ‘male’, and 1 preferred not to answer. Participants were assigned to either the Likert or the Choice condition, and saw items in only one of the two measurement techniques.
Each item in Likert was split into four sub-items: . Similarly, each Choice item was split into two sub-items: . A total of 16 participants provided Likert-scale judgments, whereas 14 provided Choice responses to the items. Most Likert sub-items () received at least 4-5 judgments, with all items receiving at least 3 judgments. The average no. of ratings per sub-item were 4. All Choice sub-items received 7 judgments per pair. Participants never saw more than one sub-item of the same item (i.e., participants couldn’t rate both and in the Likert study).
Main study
To obtain reliable ratings across the full set of EWoK-core-1.0 items, we collected at least 5 responses per item. To control for quality of data, we excluded 59 participants whose inter-subject correlation (compared with average ratings on items from other participant who rated the same items as this subject) was below . Each EWoK-core-1.0 generated item (corresponding to 5 variable-populated versions) was split into four subparts: and presented in a Likert setting using the same prompt as that used for LLMs (Appendix A.4; adapted to presentation in a web browser with a free-form text-box for human input). Items were presented so that each participant only saw one of 5 variable-populated variants and one of the four possible sub-items . Therefore, participants provided sensibility judgments independently of any other subparts of the item, closely matching the conditions for LLM evaluation. Each participant saw a list of 57 items on average, depending on the domain (188 for agent properties and 27 for material properties).
A.2 Evaluated models
The full set of evaluated models are as follows: gpt2-xl (Radford et al., 2019), phi-1 (Gunasekar et al., 2023), phi-1.5, phi-2 (Li et al., 2023), gemma-2b, gemma-1.1-2b-it, gemma-7b, gemma-1.1-7b-it (Gemma et al., 2024), mpt-7b, mpt-7b-chat, mpt-30b, mpt-30b-chat (MosaicML, 2023), falcon-7b, falcon-7b-instruct, falcon-40b, falcon-40b-instruct (Almazrouei et al., 2023), mistral-7b-v0.1 (Jiang et al., 2023), mixtral-8x7b-v0.1 (Jiang et al., 2024), Meta-Llama-3-8B, and Meta-Llama-3-70B (AI@Meta, 2024). All models were accessed via HuggingFace transformers (Wolf et al., 2019), and all experiments were run on a x GB GPU cluster.
A.3 Text completion experiments
For the prompting-based evaluations, Choice and Likert, we support two different generation options: free and constrained. For free generation, the LLM may greedily sample up to 20 tokens, and we match the first occurrence of a valid response (a numeral between 1–2 or 1–5) with a regular expression. Such a strategy avoids penalizing completions that begin with text or white-space, but doesn’t guide the model to produce a valid response. For constrained generation, the LLM may greedily sample from a restricted set of tokens, either 1–2 or 1–5, constrained using logit masking (Willard and Louf, 2023). Such a strategy enforces well-structured responses, but requires a restricted response format. In addition to variation in generation options, we support both zero- and few-shot prompting. In Figure 6, the prompting results we report use -shot constrained generation, as it yields the highest performance among our space of tested strategies.
A.4 Prompts
For the two prompt-based evaluations, Choice and Likert, we include below our exact prompt templates. The Likert prompt was additionally used verbatim for human data evaluation.
Choice Template:
Likert Template:
A.5 Mixed effects modeling
To evaluate joint effects of domains, item design factors, and item surface features on LLM performance, we entered those predictors into a mixed effects logistic regression model implemented in R using lme4. The model formula is: Accuracy 0 + Domain + ContextContrast + TargetContrast + ContextType + Frequency + NumWords + (1|Model) + (1|Item)
See Section 3 for possible values for Domain, ContextType, ContextContrast, and targetContrast. Domain effects were estimated relative to a 0 intercept. ContextType, ContextContrast and TargetContrast had deviation contrast coding. Relative word frequency and number of words per item were computed as (for either or ); these values were z-scored before being entered as regressors. Model refers to an LLM being used, and Item refers to each individual item, i.e. a minimal pair of pairs. The model was fit on item-level binary accuracy data, with 1 row per target sentence. The results we report are from model performance using LogProbs evaluation type. See Table B.2 for results.
Appendix B Results
B.1 Human study
Pilot study: Human judgments are invariant to Likert or Choice measurement
We determined humans are closely aligned when providing Likert-scale or Choice judgments (Fig. 7). Likert-scale judgments are less dependent on the specific set up (making a Choicejudgment requires a specific framing eliciting a comparison of two contexts given a target). In order to have data allowing more flexible comparisons we decided to stick to Likert-scale judgments for the full data collection in the main study.
B.2 Model evaluation
[ ] colspec=Q[]Q[]Q[],
Model Mean LogProbs Accuracy Range
human 0.951 0.942-0.957
gpt2_xl 0.655 0.645-0.662
phi_1 0.522 0.517-0.530
phi_1_5 0.727 0.686-0.756
phi_2 0.718 0.696-0.771
gemma_2b 0.678 0.657-0.705
gemma_7b 0.714 0.697-0.734
gemma_1.1_2b 0.654 0.643-0.673
gemma_1.1_7b 0.720 0.708-0.736
mpt_7b 0.733 0.716-0.745
mpt_7b_chat 0.751 0.73-0.769
mpt_30b 0.757 0.743-0.786
mpt_30b_chat 0.771 0.758-0.777
falcon_7b 0.723 0.713-0.744
falcon_7b_instruct 0.717 0.701-0.730
falcon_40b 0.783 0.775-0.789
falcon_40b_instruct 0.801 0.790-0.810
Mistral_7B 0.775 0.768-0.783
Mixtral_8x7B 0.784 0.774-0.794
Llama_3_8B 0.746 0.740-0.756
Llama_3_70B 0.775 0.759-0.787
[ ] colspec=Q[]Q[]Q[]Q[],
Domain LLM (average) LLM (best) Human
social
interactions 0.859 0.945 1.000
social
properties 0.839 0.905 0.997
material
dynamics 0.816 0.885 0.911
social
relations 0.761 0.856 0.992
quantitative
properties 0.725 0.823 0.986
physical
dynamics 0.706 0.920 0.833
agent
properties 0.683 0.778 0.975
physical
interactions 0.672 0.759 0.910
material
properties 0.669 0.755 0.921
physical
relations 0.627 0.723 0.886
spatial
relations 0.615 0.749 0.958
[ caption=Domain, design, and surface level features jointly contribute to LLM performance. *; **; *** ,
] colspec=Q[]Q[]Q[],
Predictor Type Predictor Effect
domain social interactions 1.91 ***
social properties 1.79 ***
material dynamics 2.23 ***
social relations 1.27 ***
quantitative properties 1.09 ***
physical dynamics 0.88 **
agent properties 0.58 **
physical interactions 0.83 ***
material properties 0.75 **
physical relations 0.38
spatial relations 0.41
context contrast antonym vs. rest 0.09 ***
negation vs. rest 0.1 **
variable swap vs. rest 0.0
target contrast variable vs. concept swap 0.0
context type direct vs. indirect 0.2 ***
surface features word frequency 0.07 ***
number of words **
[ caption=LLM Likert accuracy on EWoK-core-1.0 with stricter inequality metric. ,
] colspec=Q[]Q[]Q[],
Model Mean Likert Acc Range
mpt_7b 0.021 0.018-0.025
mpt_7b_chat 0.100 0.089-0.113
mpt_30b 0.310 0.307-0.316
mpt_30b_chat 0.307 0.25-0.332
falcon_7b 0.003 0.003-0.003
falcon_7b_instruct 0.005 0.004-0.008
falcon_40b 0.216 0.185-0.24
falcon_40b_instruct 0.353 0.337-0.368
Mistral_7B 0.396 0.375-0.429
Mixtral_8x7B 0.511 0.471-0.533
Llama_3_8B 0.195 0.145-0.222
Llama_3_70B .588 .576-.603

### v2

@@
 Abstract
-The ability to build and leverage world models is essential for a general-purpose AI agent.
-Testing such capabilities is hard, in part because the building blocks of world models are ill-defined.
-We present Elements of World Knowledge (EWoK), a framework for evaluating world modeling in language models by testing their ability to use knowledge of a concept to match a target text with a plausible/implausible context.
-EWoK targets specific concepts from multiple knowledge domains known to be vital for world modeling in humans.
-Domains range from social interactions (help/hinder) to spatial relations (left/right).
-Both, contexts and targets are minimal pairs.
-Objects, agents, and locations in the items can be flexibly filled in enabling easy generation of multiple controlled datasets.
+The ability to build and reason about models of the world is essential for situated language understanding.
+But evaluating world modeling capabilities in modern AI systems—especially those based on language models—has proven challenging, in large part because of the difficulty of disentangling conceptual knowledge about the world from knowledge of surface co-occurrence statistics.
+This paper presents Elements of World Knowledge (EWoK), a framework for evaluating language models’ understanding of the conceptual knowledge underlying world modeling.
+EWoK targets specific concepts from multiple knowledge domains known to be important for world modeling in humans, from social interactions (help, deceive) to spatial relations (left, right).
+Objects, agents, and locations in the items can be flexibly filled in, enabling easy generation of multiple controlled datasets.
 We then introduce EWoK-core-1.0, a dataset of 4,374 items covering world knowledge domains.
-We evaluate open-weights large language models (B–B parameters) across a battery of evaluation paradigms along with a human norming study comprising 12,480 measurements.
-The overall performance of all tested models is worse than human performance, with results varying drastically across domains.
-These data highlight simple cases where even large models fail and present rich avenues for targeted research on LLM world modeling capabilities.
+We evaluate open-weights large language models (B–B parameters) and compare them with human performance.
+All tested models perform worse than humans, with results varying drastically across domains.
+Performance on social interactions and social properties was highest and performance on physical relations and spatial relations was lowest.
+Overall, this dataset highlights simple cases where even large models struggle and presents rich avenues for targeted research on LLM world modeling capabilities.
 1 Introduction
-Large language models (LLMs) acquire a substantial amount of knowledge from their training data, both through direct memorization and through learning co-occurrence-based text patterns (Bender and Koller, 2020; Grand et al., 2022; Pavlick, 2022).
+Large language models (LLMs) acquire a substantial amount of knowledge from their training data (Bender and Koller, 2020; Grand et al., 2022; Pavlick, 2022).
 This knowledge comprises both knowledge about language (e.g.
@@
 social conventions and physical properties of objects).
-Contemporary LLMs’ knowledge of language is very robust (Mahowald et al., 2024), as evidenced by the fluency of the text that they generate.
+Contemporary LLMs demonstrate substantial knowledge of language (Mahowald et al., 2024), as evidenced by the fluency of the text that they generate.
 But how robust is their understanding of the basic social, physical, and relational concepts that are foundational to our everyday experience?
-In this paper, we present Elements of World Knowledge (EWoK)111Data and associated code are available at: http://ewok-core.github.io, a flexible framework for evaluating world modeling in LLMs (see Figure 1 for an overview).
-The EWoK framework consists of: (a) several domains that constitute the foundation for basic human world knowledge and are processed by dedicated cognitive and neural systems; (b) specific concepts within each domain; (c) a set of item templates (and modular components to procedurally generate these templates) that test knowledge of these concepts by contrasting plausible and implausible context–target combinations; (d) a set of fillers to populate the templates, such that each template can be used multiple times; (e) a pipeline to generate a specific set of items (a dataset) based on these source materials.
+In this paper, we present Elements of World Knowledge (EWoK)111Data and associated code can be accessed at http://ewok-core.github.io , a flexible framework for evaluating world modeling in LLMs (Figure 1).
+The EWoK framework consists of: (a) several domains that constitute the foundation for basic human world knowledge and are processed by dedicated cognitive and neural systems; (b) a set of concepts within each domain; (c) a set of templates (and modular components to procedurally generate these templates) that test knowledge of specific concepts by contrasting plausible and implausible context–target combinations; (d) a set of fillers to populate the templates, such that each template can be used multiple times; (e) a pipeline to generate a specific set of items (a dataset) based on these source materials; and (f) an evaluation procedure that measures language models’ ability to distinguish plausible and implausible items.
 Why elements?
-Our framework targets specific cognitive concepts (or concept pairs, such as friend/enemy).
-Concept knowledge is not limited to definitions, but can be used productively across a wide range of scenarios.
-Thus concepts leveraged in context are the first-class object of the EWoK framework, as opposed to individual sentences or facts.
-This approach stands in contrast to many NLP benchmarks, which often aim to evaluate knowledge based on individual items, or to develop sophisticated tasks that require drawing on many concepts simultaneously.
-Although real-world tasks can be more useful for comparing model performance, individual item complexity makes it hard to assess why a model fails.
-For instance, LLMs exhibit mixed performance on theory of mind tasks, which has led to disagreements about whether LLMs fail at inferring mental states or auxiliary cognitive abilities, such as knowing what it means for a box to be “transparent” (Kosinski, 2023; Ullman, 2023).
-Our framework mitigates this problem by explicitly linking the items with the concepts that they test.
+Our framework targets specific cognitive concepts (or concept pairs, such as left/right).
+Concept knowledge is not limited to memorized definitions, but can be used productively across a wide range of scenarios.
+Thus concepts (not isolated sentences or facts) are the primary focus of the EWoK framework.
+This approach stands in contrast to many NLP datasets, which often include naturalistic, lengthy stimuli that draw on multiple unspecified concepts.
+Although naturalistic tasks are useful for evaluating model performance in real-life settings, individual item complexity and lack of controls can make it difficult to assess why a model fails.
+For instance, LLMs exhibit mixed performance on theory of mind tasks, which has led to disagreements about whether LLMs fail at inferring mental states or auxiliary cognitive abilities, such as knowing what it means for a box to be “transparent” (Kosinski, 2024; Ullman, 2023).
+Our framework mitigates this problem by explicitly linking individual items with the specific concepts that they test.
 Why cognition-inspired?
 World knowledge is a notoriously fuzzy capability; which knowledge domains should one focus on?
-Here, we select a range of domains that have been shown to recruit dedicated cognitive and/or neural machinery in humans, such as knowledge of intuitive physics (McCloskey, 1983; Battaglia et al., 2013), knowledge of physical and spatial relations (Hafri and Firestone, 2021), intuitive number sense (Dehaene, 2011), social reasoning (Carlson et al., 2013), and reasoning about agents that involves both physical and social knowledge (Liu et al., 2024).
+Dataset design in NLP is often centered around data availability, with examples sourced from video captions, Wikipedia articles, or social media logs.
+In contrast, here, we systematically select a range of knowledge domains that have been shown to recruit dedicated cognitive and/or neural machinery in humans, such as knowledge of intuitive physics (McCloskey, 1983; Battaglia et al., 2013), knowledge of physical and spatial relations (Hafri and Firestone, 2021), intuitive number sense (Dehaene, 2011), social reasoning (Carlson et al., 2013; Thomas, 2024), and reasoning about agents that involves both physical and social knowledge (Liu et al., 2024).
 These knowledge domains are not specific to language (Jackendoff, 2002); in fact, many are present in preverbal infants (Spelke and Kinzler, 2007).
-However, language contains a rich amount of information that reflects grounded world knowledge (Roads and Love, 2020; Abdou et al., 2021; Patel and Pavlick, 2021), such that LLMs might in principle acquire such domain-specific knowledge from text alone.
+At the same time, traces of these signals are deeply embedded in our daily communicated language.
+Text itself thus contains rich information that reflects grounded world knowledge (Roads and Love, 2020; Abdou et al., 2021; Patel and Pavlick, 2021), and it may be reasonable to predict that LLMs would acquire this domain-specific knowledge from text alone.
 Why plausibility?
-To evaluate basic world modeling, we use combinations of plausible vs.
+Concept understanding is often fuzzy.
+Thus, deciding whether a factual statement is true or false in isolation is often ill-defined.
+Instead, we use combinations of plausible vs.
 implausible context–target pairs.
-Plausibility here serves as a proxy for factual accuracy: instead of deciding whether a factual statement is true or false, a model needs to determine whether a given scenario makes sense (is plausible).
-Having an accurate world model is necessary for consistently distinguishing plausible and implausible scenarios no matter how they are worded.
+In EWoK, a model needs to determine which scenarios make more sense (more plausible).
+Having an accurate world model would enable a model to consistently distinguish plausible and implausible scenarios no matter how they are worded.
 Why minimal pairs (of pairs)?
-Both contexts and targets in EWoK have a minimal-pairs design, such that a specific targeted change to a sentence (e.g., left right) results in an opposite result (plausible/implausible combination).
+Both contexts and targets in EWoK have a minimal-pairs design, such that a specific targeted change to a sentence (e.g., left right) yields an opposite effect (e.g., plausible implausible).
 This approach can help identify specific manipulations that LLMs are and are not sensitive to, with the goal of targeted diagnostics.
-It can also be leveraged in the future for mechanistic interpretability research.
+Such controlled manipulations are also particularly well-suited for mechanistic interpretability research, which may seek to explain the circuits through which models deploy world knowledge in context.
 Why context–target combinations?
 LLMs have a remarkable capacity for memorization, such that many plausible and implausible sentences can be distinguished solely based on their presence in the training data (e.g., The fox chased the rabbit is more common than The rabbit chased the fox).
-In contrast, our framework tests LLMs’ ability to evaluate contextual plausibility, such that the same exact target (The piano is left of Ali.) is either plausible or implausible depending on the context (see Figure 1 right).
+In contrast, our framework tests LLMs’ ability to evaluate contextual plausibility, such that the same exact target (The piano is left of Ali) is either plausible or implausible depending on the context (see Figure 1 right).
+Establishing this control is critical for isolating contextually-sensitive world knowledge from heuristic knowledge associated with surface forms.
 Our paper is structured as follows.
-In Section 2, we review prior research related to this topic.
+In Section 2, we review prior research related to the topics explored in this work.
 In Section 3, we describe the core components of the EWoK framework.
-In Sections 4 and 5, we describe our evaluation strategy and the practical steps that we take to reduce the risk of dataset contamination.
+In Section 4, we describe our evaluation strategy.
 In Section 5, we show LLM performance on EWoK-core-1.0, a dataset generated via the EWoK framework.
 In Section 6, we discuss our results and future prospects for this line of work.
+Finally, in Section 7, we discuss dataset release considerations.
+| Domain | Templates | Concepts | Example | ||
+| Social Interactions | 165 | 16 |
+| Social Properties | 185 | 16 |
+| Social Relations | 785 | 15 |
+| Physical Interactions | 280 | 20 |
+| Physical Dynamics | 75 | 12 |
+| Physical Relations | 435 | 20 |
+| Material Dynamics | 780 | 21 |
+| Material Properties | 125 | 18 |
+| Agent Properties | 1130 | 22 |
+| Quantitative Properties | 195 | 18 |
+| Spatial Relations | 245 | 14 |
 2 Related Work
@@
 LLMs often struggle on such commonsense benchmarks, likely due to the reporting bias in their training data (Shwartz and Choi, 2020): conversations and texts typically do not include commonly observed or obvious information (Gordon and Van Durme, 2013).
-Thus, although our items target basic world knowledge and are not designed to be challenging, LLMs might still struggle with them due to the reporting bias.
+Thus, although our items target basic world knowledge and are not designed to be challenging, it is possible that LLMs could still struggle with them due to the reporting bias.
 One specific version of reporting bias affects perceptually grounded knowledge.
@@
 The recognizing textual entailment (RTE) task Dagan et al.
-(2010) poses two sentences (a text expression and and hypothesis ) to a system and asks it to determine whether follows from .
+(2010) poses two sentences to a system (a text expression and hypothesis ) and asks that it determine whether follows from .
 The natural language inference (NLI) task follows a similar challenge (Bowman et al., 2015; Williams et al., 2018; Conneau et al., 2018) and involves making a 3-way judgment about whether a premise entails, contradicts, or is neutral relative to a hypothesis.
@@
 Though widely successful as a challenge, a large body of subsequent work has highlighted issues with RTE- and NLI-style evaluation: language models can often use heuristics (such as artifacts left behind by human annotators and lexical statistics) to “solve” the task without meaningful semantic understanding or reasoning (Poliak et al., 2018; Liu et al., 2020; McCoy et al., 2020; Gururangan et al., 2018).
-We address this limitation by (1) posing the task as a minimal pair where each of two targets must be preferred over its alternative given the right context, making reliance on target plausibility alone impossible; (2) annotating minimal pair contrast type to test whether which item design features drive model performance; and (3) testing the relationship between LLM performance and surface-level item properties, such as item length, average word frequency, and performance of a baseline bag-of-words embedding model.
-Our approach to dataset design is similar in spirit to the bAbi framework (Weston et al., 2016), which used simple synthetic tasks probing world knowledge and reasoning; however, our items are both simpler in design (they target individual concepts and do not, as of yet, require multi-chain reasoning or transitive inference) and harder in practice (a minimal pair, context-dependent design greatly reduces the availability of response heuristics, a serious problem in bAbi; Kaushik and Lipton, 2018).
+We address this limitation by (1) posing the task as a minimal pair, where each of two targets is held static while paired contexts are used to modulate preferences, rendering it impossible to rely solely on target plausibility; (2) annotating minimal pair contrast type to test whether any item design feature drives model performance; and (3) testing the relationship between LLM performance and surface-level item properties, such as item length, average word frequency, and performance of a baseline bag-of-words embedding model.
+Our approach to dataset design is similar in spirit to the bAbi framework (Weston et al., 2016), which used simple synthetic tasks probing world knowledge and reasoning; however, our items are both simpler in design (they target individual concepts and do not require multi-chain reasoning) and are harder in practice (a minimal pair, context-dependent design greatly reduces the availability of response heuristics, a serious problem in bAbi; Kaushik and Lipton, 2018).
 The minimal pair design is common in datasets inspired by psycholinguistics and cognitive science, such as SyntaxGym (Gauthier et al., 2020), BLiMP (Warstadt et al., 2020), and COMPS (Misra et al., 2023).
-In particular, it has previously been used to evaluate the models’ ability to distinguish plausible and implausible events (Pedinotti et al., 2021; Kauf et al., 2023), a task that draws heavily on commonsense knowledge.
+In particular, it has previously been used to test models’ ability to distinguish plausible and implausible events (Pedinotti et al., 2021; Kauf et al., 2023), a task that draws heavily on commonsense knowledge.
 The popular Winograd Schema Challenge also had a minimal pairs setup (Levesque et al., 2012), even though that requirement was relaxed in later versions.
 We here extend this approach by employing a minimal pairs-of-pairs design, where both context and target sentences have a minimal pair counterpart.
-Until 2023, the dominant approach of assessing language model performance on a minimal pair has been to calculate each item’s (pseudo) log probability under the model.
+Before prompt-based evaluation became commonplace, the dominant approach for assessing language model performance on a minimal pair has been to calculate each item’s (pseudo) log probability under the model.
 This method is effective at distinguishing grammatical and ungrammatical sentences (e.g., Warstadt et al., 2020), plausible and implausible events (Kauf et al., 2023), and relevant vs.
@@
 Yet raw log probabilities reflect a number of surface-level properties of the input, such as word frequency (Kauf et al., 2023) and the number of possible paraphrases (Holtzman et al., 2021).
-An alternative approach, recently made possible with more powerful LLMs, is to prompt an LLM to to rate item plausibility, either absolute (on a Likert scale) or relative to the other item in the minimal pair.
+An alternative approach, recently made possible with more powerful LLMs, is to prompt an LLM to rate item plausibility, either absolute (on a Likert scale) or relative to the other item in the minimal pair.
 This approach can theoretically result in more task-specific estimates.
@@
 3 The Framework
-We provide a flexible generative synthetic data pipeline (Figure 1), capable of producing many diverse datasets, each with unique specifications and statistics, while preserving metadata and decision traces.
-In Section 6, we use this framework to generate EWoK-core-1.0, a systematic, broad-coverage, context-sensitive world knowledge dataset containing 4,400 items.
-Item format
-Effective use of world knowledge incorporates both access to a rich set of priors about the world’s structure and the ability to integrate this information on-the-fly with surrounding context.
+We provide a flexible generative synthetic data pipeline, capable of producing many diverse datasets, each with unique specifications and statistics, while preserving metadata and decision traces.
+Here, we discuss the design principles that make it possible to reuse the EWoK framework for systematically generating various world knowledge datasets.
+In Section 5, we use this framework to generate EWoK-core-1.0, a systematic, broad-coverage, context-sensitive world knowledge dataset containing items that probe concepts (for a full concept list, see Table A3).
+Items test concept knowledge in context.
+Effective use of world knowledge incorporates both access to a rich set of priors about the world’s structure and the ability to integrate this knowledge on-the-fly with information about the current environment.
 Thus, our items challenge models to leverage concepts in context.
@@
 Thus, base target probabilities and cannot serve as plausibility cues: a model has to rely on context to establish plausibility.
-Domains & concepts
-EWoK is designed around domains of general world knowledge, with the current set shown in Figure 3.
-Each domain includes a set of concepts.
-For example, the domain social relations includes friend, enemy, teacher, student, boss and subordinate, and so on.
-Dataset generation procedure
+Concepts comprise domains of world knowledge.
+EWoK is designed around domains of general world knowledge, with the current set of 11 domains shown in Table 1.
+We selected a range of domains that have been shown to recruit dedicated cognitive and/or neural machinery in humans, backed by extensive literature in cognitive science (introduced in Sec.
+1, “Why cognition inspired?”).
+Domains were contributed based on past literature during the development of EWoK by a team of experts in the field (authors of this paper: professors, post-docs, and grad students in cognitive science and neuroscience across a few institutions).
+Each domain includes a set of concepts, also curated by the team.
+For example, the domain social relations includes friend, enemy, teacher, student, boss, subordinate, and others.
+The number of concepts in each domain is mentioned in Table 1.
+Items are generated from concept-specific templates.
 Each concept is associated with several items that test knowledge of the concept (often, but not always, by contrasting it with another concept).
-Items are created in a flexible yet controlled manner using the EWoK dataset generation procedure.
-At the core of the framework, we have atomic units and combination rules that, subject to constraints, lead to the generation of templates which are then populated with fillers—this leads to generating many more carefully-controlled items than a single-step template-filling approach could generate.
-Contexts & targets
-A target is a simple sentence that incorporates a concept; each domain has one or more associated targets depending on the lexical properties of the concepts in that domain.
-A contrasting target pair is generated using one of two mechanisms: concept swap, which contrasts the same target with different concepts filled in, and variable swap, which swaps two objects or agents mentioned in the target (only possible for certain concepts).
-For instance, {agent1} is to the left of {agent2} can be contrasted with {agent1} is to the right of {agent2} (concept swap) or with {agent2} is to the left of {agent1} (variable swap).
-A context pair is one or more minimal pair(s) of sentences that can each be paired with a target pair such that but not matches and but not matches .
-They are typically specific to an opposing concept pair (left/right) or to a single concept (left; only possible when variable swap is allowed).
-Similarly, a contrasting context pair is generated using one of two mechanisms, namely, a filler swap and a variable swap.
-A filler swap uses contrasting fillers to generate two versions of the item.
-A variable swap changes positions of two entities of the same kind to change the nature of the relationship or dynamic between them as long as the concept supports it.
-Templates & fillers
-Each collection of concepts, contexts, and targets can then be compiled to a set of templates, partial items with typed variables describing the range of fillers for which the intended meaning holds.
-For example, in the template {object2:can_bounce=True} bounced off {object1} from below, any object may fill {object1}, e.g., the desk or the crate, whereas {object2:can_bounce=True} must be filled by an object marked with a flag for can_bounce=True, e.g., the ball or the tire.
+Items are generated from templates, which contain placeholders for specific objects, agents, and locations.
+The placeholder value should not affect the resulting plausibility judgments; thus, a template can be populated with an arbitrarily large number of fillers, which enables generating many carefully controlled items.
+Templates consist of contrasting context and target pairs.
+A template includes 2 targets and 2 contexts.
+A target is a simple sentence that incorporates a concept.
+A context is a short sequence of words that can be paired with a target to yield either a plausible or an implausible combination.
+Context and target pairs are designed in such a way that but not matches and but not matches .
+The differences between the two targets or the two contexts result in specific types of contrasts, shown in Figure 2.
+A contrasting target pair is generated using one of two mechanisms: concept swap, which contrasts the same target with different concepts filled in, and variable swap, which swaps two objects or agents (only possible for certain targets).
+For instance, AGENT-1 is AGENT-2’s teacher can be contrasted with AGENT-1 is AGENT-2’s student (concept swap) or with AGENT-2 is AGENT-1’s teacher (variable swap).
+A contrasting context pair is generated using antonyms, negation, or variable swap.
+Finally, the templates themselves test concept knowledge in either direct or indirect way.
+Direct tests rely on immediate interpretation of a concept, whereas indirect tests leverage the probabilistic relationship between concept and target ( doesn’t have to be true given but it’s more likely than given ).
+Templates enable flexible yet controlled generation of items using typed fillers.
+Placeholders in templates (e.g., OBJECT or AGENT) can be restricted to only allow fillers of specific types.
+For example, in the template {object2:can_bounce=True} bounced off {object1} from below, {object2:can_bounce=True} must be filled by an object marked with a flag for can_bounce=True, e.g., the ball.
+These type restrictions prevent generation of semantically anomalous or incomprehensible sentences (e.g., The blanket bounced off the table from below).
+Type restrictions were specified by the author team during the template development stage and later validated via human behavioral ratings (Section 4.4 and Appendix A.2).
+To create an initial set of object and location fillers, we generated a set of arbitrary fillers based on constraints specified in templates (e.g., objects that can bounce, that have a smell) that was filtered down to ensure the ‘naturalness’ of the items that include these fillers.
+The agent fillers were selected from the list of culturally diverse high frequency male, female and non-gendered baby names.
 Across domains we list over filler items across classes with type restrictions.
-When populating the templates with fillers, users can specify various custom parameters, e.g.: (1) specify the number of items to generate from each template; each full set of items is referred to as a version; (2) select whether fillers should be held constant across all items in a version or allowed to vary; and (3) apply transformations to filler restrictions at compile-time, for instance restricting all agents to take non-Western names via agent->agent:western=False, or swapping all objects with nonce words via object->nonword.
+The full set of fillers is provided in the paper repo.
+When populating the templates with fillers, users wanting to use the EWoK framework to construct their own instance of an EWoK-core-1.0-like dataset can specify several parameters, e.g.: (1) the number of items to generate from each template; each full set of items is referred to as a version (corresponding to a random seed); (2) whether fillers should be held constant across all items in a version or allowed to vary; (3) whether to apply transformations to filler restrictions at compile-time, for instance restricting all agents to take non-Western names via agent->agent:western=False, or swapping all objects with nonce-words via object->nonword.
 Such flexibility allows for controlled experimentation of the features modulating model performance.
-With these arguments and others easily accessible from the command line interface, users are supported to generate an enormous space of EWoK variations, and may easily add new domains, concepts, fillers, and more.
+With these arguments and others easily accessible from the command line interface, users are supported to generate many EWoK variations, and can additionally extend EWoK to add new domains, concepts, and fillers.
 4 Evaluation
-Using the EWoK framework, we compiled the EWoK-core-1.0 dataset by generating unique fixed substitutions of filler items across templates from domains.
-In this section, we describe the evaluation of EWoK-core-1.0 via three different evaluation paradigms: traditional plausibility estimates via querying the probability of sentences under the model, LogProbs, as well as two prompt-based strategies, Likert and Choice.
+Using the EWoK framework, we compiled the EWoK-core-1.0 dataset.
+To do so, our team (which includes experts in specific domains of cognitive science) curated a list of templates that test knowledge of concepts from domains.
+These templates were then populated with fillers that were randomly sampled from our initial filler set (described in Section 3), resulting in dataset versions.
+This strategy enables us to explicitly measure the variability associated with the random sampling of filler items (which, in principle, should not affect the results).
+For EWoK-core-1.0, we chose to hold the set of fillers constant across all items within each version, although that constraint can be relaxed if needed (see Appendix A.1).
+We evaluated LLM performance on EWoK-core-1.0 using three different approaches: traditional plausibility estimates via querying the probability of sentences under the model, LogProbs, as well as two prompt-based strategies, Likert and Choice.
 The majority of the results reported use the LogProbs evaluation method, which allows comparing a wide range of base and finetuned models.
@@
 For the prompt-based evaluations, we collected data from both LLMs as well as human participants using paired identical prompts.
-Drawing inspiration from comparative psychology, such an approach of matched evaluation has been proposed as a way to support increasingly fair evaluations of LLMs, allowing for more direct comparison of performance, consistency, and context-sensitivity with human participants (Lampinen, 2022).
+Drawing inspiration from comparative psychology, such an approach of matched evaluation has been proposed as a way to support increasingly fair evaluations of LLMs, allowing for more direct comparison of performance, consistency, and context-sensitivity with human participants (Lampinen, 2024).
 4.1 Scoring metrics
 For LogProbs evaluation, we use token-level LLM probabilities to calculate as a sum of conditional log probabilities of each token: , where is the vector of tokens composing the target .
-For Likert, participants (humans and models) are prompted to rate the plausibility of the concatenation of each and pair on a 1–5 scale.
+For Likert, participants (humans and models) are prompted to rate the plausibility of the concatenation of each and pair on a 1–5 scale222A 1–5 point scale was chosen to distinguish between degrees of plausibility as opposed to solely absolute direction.
+Several items in our dataset were distinguishable through this increased granularity, e.g., a 1 vs 2 or 4 vs 5 that would not have been distinguishable otherwise..
+This is functionally similar to the LogProbs condition, but explicitly based on the model’s generative behavior as opposed to intrinsic to its scoring.
 For Choice, participants (humans and models) are presented with and , followed by a single target ( or ), and then prompted to select the context ( or ) that better matches the target.
-For Likert and Choice, details about text generation hyperparameters can be found in A.3 and exact prompt templates can be found in Appendix A.4.
+This is another prompting approach that uses the language model discriminatively with access to both items.
+For Likert and Choice, details about text generation hyperparameters can be found in A.4 and exact prompt templates can be found in Appendix A.5.
 The metric for correctness of a given item is the recovery of the designed item structure that and , where score reflects for LogProbs, the integer rating for Likert, and correct context index selection for Choice.
-In all cases, participants must correctly identify both matches to get the full score ( point).
+In all cases, models must correctly identify both matches to get the full score ( point).
 Identification of only one match receives points.
-In the case of the Likert task, if a model returns the same rating for both pairs, the model receives points.
+In the case of the Likert task, which has a coarser integer scale, if a model returns the same rating for both pairs, the model receives points.
 Such a paradigm supports a trivial baseline for all scenarios.
-Either a random coin flip or a deterministic model generating the same response for each query independent of context will trivially achieve this baseline.
+Even a random coin flip or a deterministic model generating the same response for each query independent of context will trivially achieve this baseline.
+For prompt-based evaluations (Likert and Choice), we used the same prompting setup across all models to assess them in the same way as humans; however, unlike humans, models were additionally provided with -shot examples.
+These examples did not come from any of our domains and were solely meant to outline the required response formatting including one positive and one negative example.
+See Appendix A.5 for more details.
+Our initial experiments revealed that supplying these examples substantially improves model performance, providing us with an opportunity to evaluate their “best shot” at EWoK-core-1.0.
 4.2 Models
@@
 Since we evaluate LLMs and humans on identical prompts, it was critical to have complete flexibility in formatting.
-The full set of evaluated models are listed in Fig.
+The full set of evaluated LLMs are listed in Fig.
 3 as well as in Appendix A.3.
-4.3 Surface-level item properties
 We additionally tested a baseline bag-of-words model based on word2vec embeddings (Mikolov et al., 2013).
@@
 The context/target match was determined by a cosine similarity metric, such that an item is scored correctly if > (and vice versa for ).
-We then examined the extent to which LLM and human performance correlates with the word2vec baseline.
-We also tested whether LLM performance correlates with the number of words in an item, as well as with average word frequency in an item using unigram counts from the Google Ngrams (Michel et al., 2011) 2012 American English corpus.
+Word-embedding-based models provide useful baselines, and there is a long tradition of work in NLI and related tasks utilizing such easy-to-understand baselines to assess task difficulty and the presence of alternative solution strategies Naik et al.
+(2018).
+Past NLI work has inadvertently fallen prey to design biases, making it easy for models to rely on heuristics, such as word overlap (McCoy et al., 2020).
+Additionally, word embeddings can contain rich semantic information (Grand et al., 2022) that may help solve the task by assessing semantic relatedness between words in the context and the target.
+Using bag-of-words baselines helps understand how robust the task is in the face of such heuristics.
+4.3 Surface-level item properties
+To determine the influence of surface-level item properties on model performance, we tested whether LLM performance correlates with the number of words in an item, as well as with average word frequency in an item.
+Word frequency was determined using unigram counts from the Google Ngrams (Michel et al., 2011) 2012 American English corpus.
 4.4 Human data
 For independent norming of the items in EWoK-core-1.0, we collected data from human participants.
-We recruited a total of =1,262 participants (591 female, 579 male, 27 other; median age 36; all US-residents with English as their first language) via Prolific, an online study platform.
-Raters with poor agreement with others () were excluded.
-The task participants performed was nearly identical to the Likert version of the task for LLMs (in a pilot study, we determined that human results for Likert and Choice show correlation; see Appendix A.1 & B.1 for more details).
-5 Release Considerations
-Our release-related goals are to (a) reduce the chances of accidental incorporation of EWoK items into LLMs’ training data and (b) promote accountability and reporting when such incorporation is done intentionally.
-Thus:
-EWoK-core-1.0 is released on HuggingFace Datasets (Lhoest et al., 2021) with gated user access to prevent scrapers from accessing it automatically.
-Users will simply accept a CC-BY license and accompanying Terms of Use (TOU) wherein they will agree to explicitly report any instances when a language model was trained on the EWoK-core-1.0 items, and will be granted access automatically.
-Link: https://huggingface.co/datasets/ewok-core/EWoK-core-1.0
-The code for the EWoK item generation framework is shared in a separate repository on GitHub, with template files downloadable as password-protected archives to prevent automatic scraping.
-The repository is also protected with a TOU that requires anyone training or fine-tuning on any data generated using EWoK to report that fact.
-Link: https://github.com/ewok-core/ewok
-The code required to replicate the results in this paper, along with human study and model performance data, is shared as a separate GitHub repository following the same protections.
-Link: https://github.com/ewok-core/ewok-paper.
-Although not flawless, this strategy allows us to minimize the chances of accidental incorporation of our data into training sets, as well as to send a clear signal to model developers about the importance of accountable reporting of evaluation dataset use during model training.
+The task human participants performed was nearly identical to the Likert version of the task for LLMs (in a pilot study, we determined that human results for Likert and Choice show Pearson’s R=0.96 correlation, and decided to drop the Choice condition for the full data collection; see Appendix A.2 (experimental design) & Appendix B.1 (results) for pilot details).
+Data collection
+To obtain reliable ratings across the full set of EWoK-core-1.0 items, we collected at least 5 responses per item from a total of N=1,262 participants (591 female, 579 male, 27 other/unknown; median age 36; all US-residents who reported English as their first language).
+Participants were recruited via Prolific (https://prolific.co), an online study platform.
+To control for quality of data, we excluded 59 participants whose inter-subject Pearson correlation coefficient (compared with average ratings on items from other participant who rated the same items as this subject) was below .
+Each EWoK-core-1.0 generated item (corresponding to 5 filler-populated versions; see Sec.
+3, “Templates & fillers” and Sec.
+A.1) was split into four subparts: , , , and presented in a Likert setting using the same prompt as that used for LLMs (Appendix A.5; adapted to presentation in a web browser with a free-form text-box for human input).
+Items were presented so that each participant only saw one of 5 filler-populated variants and one of the four possible sub-items .
+Therefore, participants provided sensibility judgments independently of any other subparts of the item, closely matching the conditions for LLM evaluation.
+Each participant rated an average of 57 items (all items came from a single domain).
+The resultant average inter-subject Pearson correlation across all items was 0.744, where the inter-subject correlation was calculated per-item based on all the participants that rated that item.
+Aggregation and evaluation
+We average ratings for a given - pair across all (typically, 5) participants that rated the pair.
+To obtain human norms, we compare the averaged Likert ratings of subparts, awarding point for each half comparison, i.e., and , similar to evaluating LLMs as described in Section 4.1.
+In case of a tie in scores for a half-comparison, we award 0 points.
 5 Experiments: EWoK-core-1.0
-EWoK-core-1.0 is challenging for LLMs
-Despite the fact that EWoK-core-1.0 was not designed to be a hard dataset, we found that even the larger models generally perform much below humans: the best model tested, falcon-40b-instruct, has an accuracy of whereas humans perform at (Table AB.2).
-As expected, larger models tend to do better, although the performance gains are modest.
+EWoK-core-1.0 can be challenging for LLMs
+Although EWoK-core-1.0 was not designed to be a difficult dataset, we found that even large open models perform well below human baselines: the best model tested, falcon-40b-instruct, yields a mean accuracy of whereas mean human accuracy is (Table AB.2).
+As expected, larger models tend to do better, although it is not a sole predictor of performance.
 Instruction tuning does not consistently increase or decrease LLM performance under the LogProbs metric.
+EWoK-core-1.0 was later evaluated on closed frontier models via HELM (Liang et al., 2023) in January of 2025.
+That evaluation was performed using Choice prompting given that LogProbs for closed models are not available.
+The highest performing closed models yield performance , which is still below humans () but not by much (see Table 2).
+| Model | Mean Accuracy |
+| GPT-4 Turbo (2024-04-09) | 0.912 |
+| Claude 3.5 Sonnet (2024-06-20) | 0.911 |
+| Qwen2 Instruct (72B) | 0.901 |
+| Claude 3 Opus (2024-02-29) | 0.893 |
+| GPT-4o (2024-05-13) | 0.886 |
+| Llama 3 Instruct (70B) | 0.876 |
+| GPT-4o mini (2024-07-18) | 0.864 |
+| Claude 3 Sonnet (2024-02-29) | 0.848 |
+| Mixtral Instruct (8x22B) | 0.842 |
+| Mixtral Instruct (8x7B) | 0.837 |
+| Claude 3 Haiku (2024-03-07) | 0.829 |
+| Claude 3.5 Haiku (2024-10-22) | 0.810 |
+| GPT-3.5 Turbo (0125) | 0.797 |
+| Mistral Instruct v0.3 (7B) | 0.772 |
+| Llama 3 Instruct (8B) | 0.171 |
 Performance varies drastically by domain
 Figure 3 shows model results by domain, from easiest to hardest (see also Table AB.2).
-Social interactions is the easiest for both LLMs (mean=, best=) and humans (); spatial interactions is both hardest for LLMs (mean=, best=) and shows the largest gap with human performance (who score highly at ).
-Overall, domain difficulty is consistent across LLMs, such that hard domains are hard for all and vice versa.
+Social interactions is the easiest for both LLMs (mean=, best=) and humans (); spatial interactions is both hardest for LLMs (mean=, best=) and shows the largest performance gap with humans (who score highly at ).
+We conclude that domains that tap into social knowledge are easier for LLMs whereas domains that require relational knowledge about the physical world are hardest.
+Overall, domain difficulty is consistent across LLMs, such that the same domains are predominantly hard or easy for all LLMs.
 One notable exception is the heterogeneous performance of the phi models, with phi-1 consistently among the worst models, phi-1.5 outperforming all models and even humans on physical dynamics, and phi-2 ranging from on par with the largest models on some domains to worse than gpt2-xl on spatial relations.
@@
 In principle, these values should not affect the results ( still best matches and not ).
-However, models show somewhat different performance on the 5 different versions of the dataset, with phi-2 and phi-1.5 showing the largest performance range ( for both).
+However, models show somewhat different performance on the 5 different versions of the dataset (Table AB.2), with phi-2 and phi-1.5 showing the largest performance range ( for both).
 This heterogeneity indicates that arbitrary item choices can substantially affect model performance, such that, in our dataset and in others, “1%” improvements might not be truly meaningful because they would not generalize.
 Humans show somewhat heterogeneous performance too (range ), although that heterogeneity is driven only by a subset of the domains (Figure 3).
+One of the strengths of EWoK is that it allows us to explicitly explore this variation in details of the dataset construction process.
+In most benchmarks, these types of details in item-level construction variation are not measurable, even though these effects likely exist.
 Domain content, item design features, and surface-level item features all affect LLM performance
@@
 Figure 4 shows that all these factors affect model performance, often in different ways than they affect humans.
-Performance of a bag-of-words word2vec-based baseline on different dataset domains and versions is predictive of LLM but not human performance (although baseline performance overall is only ).
-The number of words in an item negatively affects LLM but not model performance.
+The number of words in an item negatively affects LLM but not human performance.
 Unexpectedly, we found a negative relationship between word frequency and both LLM and human performance; follow-up examination of the data showed that this effect is driven by the fact that physical-relations and spatial-relations, the two hardest domains for LLMs, have the highest word frequency.
-To evaluate relative contributions of these factors to model performance, we jointly modeled all features using mixed effects regression.
+To evaluate the relative contributions of these factors to model performance, we jointly modeled all features using mixed effects regression.
 Design features and surface-level features all contributed to model performance (Table AB.2), with word frequency having a significant positive effect and the number of words having a significant negative effect, as expected.
 Importantly, domain remained a significant predictor of performance even when accounting for other factors.
+These results indicate that surface-level factors, such as sentence length and word frequency, contribute but do not fully explain model performance.
+Finally, the baseline bag-of-words model, which selects the context with the highest word2vec cosine similarity to the target, performed much worse than LLMs, indicating that simple embedding similarity between words in contexts and targets is insufficient to explain LLM performance.
+Of note, however, is that the phi-1 model does poorly on all domains and even worse than the baseline bag-of-words model.
+We believe the overly specific nature of phi-1’s training data that consists of textbook-like demonstrations of coding problems and solutions likely makes it unable to succeed on EWoK-core-1.0.
+This helps illustrate that without exposure to a naturalistic text corpus which would contain information about the world, even a model that is fluent in following English-language instructions is not able to flexibly reason about the world.
 LogProbs yield higher accuracy than prompting
-Evaluating LLMs with LogProbs resulted in above-chance accuracy for almost all models (Figure 6).
+Evaluating LLMs with LogProbs resulted in above-chance accuracy for almost all open models (Figure 5).
 As expected, the gap between LogProbs and prompting was larger for smaller models (Hu and Frank, 2024).
-We used the same prompting setup across all models to assess them in the same way as humans (Lampinen, 2022) (although note that models are also aided via -shot examples).
-It is possible that targeted prompt-engineering or alternative evaluation strategies will result in above-LogProbs performance, a direction we leave to future work.
-Human ratings are often, but not always accurate
-Finally, we examined discrepancies between human ratings and experimental labels.
-Sometimes, the discrepancy resulted from specific fillers changing the plausibility of a pair; for instance The cooler is inside the car.
-Chao cannot see the cooler.
-is implausible because the cooler is large and the car has windows, although, for smaller objects and containers without windows, the scenario is plausible.
+While we observed this general pattern of LogProbs as the dominant strategy, it is possible that highly targeted prompt engineering may result in above-LogProbs performance.
+Human ratings are usually, but not always, accurate
+Finally, we examined discrepancies between human ratings and experimenter-designed “ground truth” labels.
+Sometimes, the discrepancy resulted from specific fillers changing the plausibility of a pair; for instance “The cooler is inside the car.
+Chao cannot see the cooler.” may be implausible because the cooler is a large object and the car has windows, although, for smaller objects and containers without windows, the scenario is more plausible.
 Sometimes, humans made mistakes.
 One such example is cardinal directions from the spatial relations domain.
-The scenario The bakery is north of Chao.
+The scenario “The bakery is north of Chao.
 Chao turns around.
-The bakery is south of Chao.
-is implausible because cardinal directions do not depend on the agent’s orientation, and yet our participants often marked it as plausible.
+The bakery is south of Chao.” is implausible because cardinal directions do not depend on the agent’s orientation, and yet our participants often marked it as plausible.
+Spatial cognition is known to be variable across individuals and cultures (Majid et al., 2004; Pitt et al., 2022), including preferential reliance on relative (‘left’/‘right’) over absolute (‘north’/‘south’) reference frames in Western cultures, which might explain why our (primarily Western) participants performed worse on items with absolute reference frames.
 Overall, human data collection is a valuable source of information on our dataset but it does not replace ground truth labels.
@@
 We present a systematic, flexible framework that can be used to test basic world knowledge in language models.
-Our goal was to develop a dataset that: (1) uses a uniform item format to probe diverse domains of physical and social knowledge, (2) contains items that probe specific concepts (“elements of world knowledge”) within these domains, (3) requires integrating information across sentences, such that the same target sentence is plausible given one context and implausible given another, and (4) consists of generic templates that can be used to generate a wide variety of items.
+Our goal was to develop a dataset that: (1) leverages a uniform item format to probe diverse domains of physical and social knowledge, (2) presents items that employ specific concepts (“elements of world knowledge”) within these domains, (3) requires integrating information across sentences such that the same target sentence is plausible given one context and implausible given another, and (4) consists of generic templates that can be used to generate a large variety of items.
 We then presented evaluation results for a set of openly available models on EWoK-core-1.0, a dataset generated using the EWoK framework.
-This dataset is moderately challenging for LLMs, with performance varying substantially across domains.
-We also show that LogProbs contain enough information for most LLMs to perform above chance and that evaluation via prompting underperforms relative to LogProbs even for better-performing models, in agreement with prior results (Hu and Frank, 2024; Hu and Levy, 2023; Hu et al., 2024; Kauf et al., 2024).
+This dataset is moderately challenging for LLMs, with performance varying substantially across domains (with social knowledge being the easiest and physical and spatial knowledge being the hardest).
+We show that LLM-generated relative item probabilities, estimated with LogProbs allow distinguishing plausible and implausible scenarios above chance for most models and domains and that evaluation via prompting underperforms relative to LogProbs even for better-performing models, in agreement with prior results (Hu and Frank, 2024; Hu and Levy, 2023; Hu et al., 2024; Kauf et al., 2024).
+Thus, we argue that a natural way to compare models is through the probabilities they output.
+This is especially the case for models before post-training/instruction-tuning, weak models, or checkpoints, all of which may not have the capability of following complex instructions (e.g., multiple choice) but have substantial knowledge that can be revealed with LogProbs.
+We discovered that our dataset has high domain distinguishability—some world knowledge domains are much easier for LLMs than others—but low model distinguishability—many LLMs perform comparably across domains.
+This pattern indicates that the specifics of the model architecture and training data are less important than the semantic content being evaluated.
+Results from all models consistently show that social knowledge is easier to learn from text alone than knowledge about the physical world.
+We also see that closed frontier models achieve high accuracy on EWoK-core-1.0.
+This is not surprising, as this dataset was not designed as a challenge benchmark but rather as a comprehensive, broad-coverage dataset of core world knowledge in humans.
+Thus, the broader framework will have value even when this specific version of the dataset gets saturated.
 The EWoK framework opens up multiple avenues for future work:
 Targeted experiments
-The flexibility of our framework allows conducting specific experiments using customized sets of fillers.
+The flexibility of our framework allows for conducting specific experiments using customized sets of fillers.
 For instance, one might investigate whether LLMs perform differently on items that include western vs.
@@
 Knowledge editing research (e.g., Meng et al., 2022, 2023) has often focused on encyclopedic knowledge; but what about knowledge of basic physical and social concepts?
-Our controlled minimal pair stimuli can allow researchers to identify and manipulate model circuits that might be selectively responsible for knowledge of specific concepts from an array of domains.
+Our controlled minimal pair stimuli can allow researchers to identify and manipulate model circuits that might be selectively responsible for knowledge of these specific concepts across several domains.
 From elements to world models
-For a model to function as a flexible and robust general purpose AI system, it needs to be able to construct, maintain, and update internal world models (Ha and Schmidhuber, 2018; LeCun, 2022) (in cognitive science, variants of such world models are also known as mental models or situation models).
+For a model to function as a flexible and robust general-purpose AI system, it needs to be able to construct, maintain, and update internal world models (Ha and Schmidhuber, 2018; LeCun, 2022) (in cognitive science, variants of such world models are also known as mental models or situation models).
 The extent to which LLMs possess and use internal world models is subject to ongoing investigation (Hao et al., 2023; Yildirim and Paul, 2024; Wong et al., 2023).
@@
 Another limitation is that we use the same prompting setup for all models.
-With tailored prompt engineering, alternative generation methods, or chain-of-thought, LLM performance could improve.
-Regardless, we expect model generation to conform with general world knowledge under a wide range of evaluation scenarios—the fact that they do not is informative.
-Finally, due to the synthetic nature of our dataset, some items are semantically weird.
-Although not necessarily a problem, an alternative is to use LLMs to populate templates.
-We leave this approach to future work noting that it can introduce confounds that inflate model performance (e.g., Panickssery et al., 2024).
+With tailored prompt engineering, alternative generation methods, or chain-of-thought reasoning, LLM performance could improve.
+Finally, due to the synthetic nature of our dataset, some items may be atypical.
+We avoid blatant semantic violations by imposing type restrictions on template variables and validating our gold labels against human ratings, so the finalized set of items in EWoK-core-1.0 are semantically valid, just perhaps uncommon.
+There may be interest in using the EWoK framework to test semantically likely, not just semantically possible items (e.g., Hu et al., 2025).
+One way to achieve this would be to use LLMs to populate templates.
+We leave this approach to future work, noting that LLM-based item generation should be used with care to avoid confounds that inflate model performance (e.g., Panickssery et al., 2024a).
+Another consequence of template-based generation is that items thus generated may not reflect the distributional statistics of natural language, and thus, log probability scores may not serve as an ideal linking function between the knowledge language models possess and the scenarios we test them on.
+Methods that explicitly identify ways to tap into LLMs’ reasoning about scenario likelihood may be necessary (e.g., identifying relevant directions in activation-space: Panickssery et al., 2024b).
+Although we can only explore a limited set of evaluations in a single paper, the EWoK framework is evaluation-method-agnostic, and promises to continue being a useful resource allowing researchers to test for world knowledge in LLMs using novel methods that may be developed in the future.
+7 Release Considerations
+Our release-related goals are to (a) reduce the chances of accidental incorporation of EWoK items into LLM training data and (b) promote accountability and reporting when such incorporation is done intentionally.
+Thus:
+EWoK-core-1.0 is released on HuggingFace Datasets (Lhoest et al., 2021) with gated user access to prevent scrapers from accessing it automatically.
+Users will simply accept a CC-BY license and accompanying Terms of Use (ToU) wherein they will agree to explicitly report any instances when a language model was trained on the EWoK-core-1.0 items, and will be granted access automatically.
+https://huggingface.co/datasets/ewok-core/EWoK-core-1.0
+The code for the EWoK item generation framework is shared in a separate repository on GitHub, with template files downloadable as password-protected archives to prevent automatic scraping.
+The repository is also protected with a ToU that requires anyone training or fine-tuning on any data generated using EWoK to report that fact.
+https://github.com/ewok-core/ewok
+The code required to replicate the results in this paper, along with human study and model performance data, is shared as a separate GitHub repository following the same protections.
+https://github.com/ewok-core/ewok-paper.
 8 Conclusion
@@
 Acknowledgements
-We thank everyone who contributed ideas on dataset design, especially Hayley Ross and Yuhan Zhang.
-We thank members of LINGO Lab at MIT for providing feedback on earlier versions of the work.
-This work was supported by the Language Mission of the MIT Quest for Intelligence.
+We thank our action editor Emiel Krahmer and the anonymous reviewers for providing valuable feedback that helped improve this work.
+We thank everyone who contributed ideas on dataset design, especially Hayley Ross and Yuhan Zhang, members of LINGO Lab at MIT for providing feedback on earlier versions of the work, and Yifan Mai for creating a HELM evaluation platform for EWoK.
+This work was supported by the Language Mission of the MIT Quest for Intelligence and by the Open Philantropy Foundation.
+A.I.
+is additionally supported by funds from Georgia Tech School of Psychology.
+J.A.
+is additionally supported by a Sloan Research Fellowship.
 Appendix A Methods
-A.1 Human data collection
+A.1 Dataset generation
+EWoK-core-1.0 is the result of 5 randomly-sampled filler-substituted versions of the templates in EWoK.
+Each version in the script below also serves as a random seed for the dataset.
+Each dataset includes 1 copy of each template, as specified by the num_fillers parameter (i.e., one unique set of fillers will be sampled for this version).
+The parameter fix_fillers makes it so that fillers are shared across templates within this version (i.e., agent1 is always filled into by Fatima, if Fatima is the first agent name sampled).
+Regardless of all of these parameters, variable constraints are always respected and additional fillers are sampled to satisfy constraints (e.g., can_bounce=true).
+| Domain | Concepts |
+| Social Interactions | help, deceive, hinder, cooperate, compete, evade, seek, chase, learn, teach, respect, insult, flirt, comfort, tease, coerce |
+| Social Properties | friendly, hostile, trustworthy, untrustworthy, tolerant, bigoted, boastful, humble, dominant, submissive, shy, extroverted, introverted, confident, warm, cold |
+| Social Relations | friend, stranger, enemy, boss, subordinate, colleague, romantic partner, teacher, student, landlord, roommate, tenant, parent, sibling, child |
+| Physical Interactions | heat, cool, lift, drop, attract, repel, throw, catch, climb on, push, pull, break, fix, collide with, revolve around, approach, kick, touch, sit on, stand on |
+| Physical Dynamics | roll, slide, fall, rise, sink, float, grow, shrink, oscillate, spin, accelerate, slow down |
+| Physical Relations | bigger, smaller, occlude, occluded, contain, inside, outside, support, supported, attached, touching, connected, block, trail, hang, tied, on, under, surround, cover |
+| Material Dynamics | fold, ripple, pour, stir, flap, splash, droop, drip, hang, pile, trickle, disperse, tap, compress, drape, squeeze, break, rip, wrinkle |
+| Material Properties | transparent, opaque, smooth, rough, heavy, light, glossy, matte, fragile, sturdy, cold, warm, soft, hard, elastic, inelastic, bouncy, not bouncy |
+| Agent Properties | see, not see, hear, not hear, smell, not smell, taste, not taste, feel, not feel, believe, doubt, imagine, prefer, have no preference, intend, not intend, make an effort, make no effort, make more effort, make less effort, choose |
+| Quantitative Properties | more, less, fewer, the same, different, a lot of, a little, none, some, all, enough, not enough, many, few, no, the most, the least, the fewest |
+| Spatial Relations | left, right, above, below, north, south, east, west, in front of, behind, close, far, toward, away |
+The full list of concepts used in EWoK-core-1.0 templates is provided in Table 3.
+A.2 Human data collection
 We collected human data in two phases: a pilot study used to validate our task on a single domain and determine the measurement technique to use for data collection (Choice or Likert), and a main study where we repeated data collection for the full set of EWoK-core-1.0 items.
@@
 Main study
-To obtain reliable ratings across the full set of EWoK-core-1.0 items, we collected at least 5 responses per item.
-To control for quality of data, we excluded 59 participants whose inter-subject correlation (compared with average ratings on items from other participant who rated the same items as this subject) was below .
-Each EWoK-core-1.0 generated item (corresponding to 5 variable-populated versions) was split into four subparts: and presented in a Likert setting using the same prompt as that used for LLMs (Appendix A.4; adapted to presentation in a web browser with a free-form text-box for human input).
-Items were presented so that each participant only saw one of 5 variable-populated variants and one of the four possible sub-items .
-Therefore, participants provided sensibility judgments independently of any other subparts of the item, closely matching the conditions for LLM evaluation.
-Each participant saw a list of 57 items on average, depending on the domain (188 for agent properties and 27 for material properties).
-A.2 Evaluated models
+To obtain reliable ratings across the full set of EWoK-core-1.0 items, we collected at least 5 responses per item from a total of N=1,262 participants (591 female, 579 male, 27 other/unknown; median age 36; all US residents who reported English as their first language).
+Details of the main study are reported in Section 4.4, therefore not repeated here.
+Data exclusion
+Some online participants may respond randomly and therefore need to be excluded.
+To evaluate the quality of individual participants’ responses, we computed Pearson R between that participant’s responses and an average of other responses for that domain.
+Participants whose correlations were <0.3 were excluded.
+Validating gold dataset labels using human data
+We used human data to validate author-assigned gold labels for items in EWoK-core-1.0.
+First, we identified all items where the majority human label did not match the gold label (e.g., the plausibility of - was higher than -).
+Then, we manually inspected those items to determine whether we could trace the discrepancy to an author error or to an error in the item generation pipeline.
+When possible, the generation pipeline was adjusted to generate the correct version of the item; alternatively, the faulty item was excluded from the dataset.
+A.3 Evaluated models
 The full set of evaluated models are as follows: gpt2-xl (Radford et al., 2019), phi-1 (Gunasekar et al., 2023), phi-1.5, phi-2 (Li et al., 2023), gemma-2b, gemma-1.1-2b-it, gemma-7b, gemma-1.1-7b-it (Gemma et al., 2024), mpt-7b, mpt-7b-chat, mpt-30b, mpt-30b-chat (MosaicML, 2023), falcon-7b, falcon-7b-instruct, falcon-40b, falcon-40b-instruct (Almazrouei et al., 2023), mistral-7b-v0.1 (Jiang et al., 2023), mixtral-8x7b-v0.1 (Jiang et al., 2024), Meta-Llama-3-8B, and Meta-Llama-3-70B (AI@Meta, 2024).
-All models were accessed via HuggingFace transformers (Wolf et al., 2019), and all experiments were run on a x GB GPU cluster.
-A.3 Text completion experiments
+All models were accessed via HuggingFace transformers (Wolf et al., 2020), and all experiments were run on a x GB GPU cluster.
+A.4 Text completion experiments
 For the prompting-based evaluations, Choice and Likert, we support two different generation options: free and constrained.
@@
 In Figure 5, the prompting results we report use -shot constrained generation, as it yields the highest performance among our space of tested strategies.
-A.4 Prompts
+A.5 Prompts
 For the two prompt-based evaluations, Choice and Likert, we include below our exact prompt templates. The Likert prompt was additionally used verbatim for human data evaluation.
@@
 Likert Template:
-A.5 Mixed effects modeling
+A.6 Mixed effects modeling
 To evaluate joint effects of domains, item design factors, and item surface features on LLM performance, we entered those predictors into a mixed effects logistic regression model implemented in R using lme4.
@@
 We determined humans are closely aligned when providing Likert-scale or Choice judgments (Fig.
-7).
+6).
 Likert-scale judgments are less dependent on the specific set up (making a Choice judgment requires a specific framing eliciting a comparison of two contexts given a target).
 In order to have data allowing more flexible comparisons we decided to stick to Likert-scale judgments for the full data collection in the main study.
+Full study
 B.2 Model evaluation
-[ ] colspec=Q[]Q[]Q[],
+[ ] colspec=>Q[]Q[]Q[],
 Model Mean LogProbs Accuracy Range
 human 0.951 0.942-0.957
+word2vec 0.542 0.539-0.547
 gpt2_xl 0.655 0.645-0.662
@@
 [ caption=LLM Likert accuracy on EWoK-core-1.0 with stricter inequality metric. ,
-] colspec=Q[]Q[]Q[],
+] colspec=>Q[]Q[]Q[]
 Model Mean Likert Acc Range
@@
 Llama_3_8B 0.195 0.145-0.222
-Llama_3_70B .588 .576-.603
+Llama_3_70B 0.588 0.576-0.603